# Measuring Casimir Force Across a Superconducting Transition

### v1

@@ Acknowledgements.
The resonance frequency of the parallel plate capacitor, consists of a suspended NbTiN membrane under the electrostatic and the Casimir interaction with the fixed NbTiN backgate, is calculation in this Supplementary Information. The schematic of the device is illustrated in Fig. S17. We focus on analyzing the fundamental mode frequency $\omega_{0}$ (or $\omega_{\rm(1,1)}$). The thin membrane is under high tension, and the height of the membrane is deviated from the equilibrium (not forced) position only by a negligible amount compared to the separation. Based on these considerations, the membrane’s dynamics are described by the following equation of motion (EoM) [98]:
$\rho t\frac{\partial^{2}H(x,y;t)}{\partial t^{2}}-\sigma t\triangledown^{2}H(x,y;t)=P_{\rm ES}(H(x,y;t))+P_{\rm Cas}(H(x,y;t))~{},$ (S.4)
where the height function $H(x,y;t)$ describes the separation between the membrane and the backgate, $\rho$ the density, $\sigma$ the tensile stress, and $t$ the thickness of the membrane. $P_{\rm ES}(H(x,y;t))$ and $P_{\rm Cas}(H(x,y;t))$ are the pressures coming from the electrostatic force and the Casimir force between the membrane and the backgate, respectively. The electrostatic and Casimir forces can be estimated using the additive Derjaguin Proximity Force Approximation (PFA) [109]. This method decomposes the NbTiN surfaces into parallel patch pairs and sums the corresponding pairwise force contributions, thus
$\rho t\frac{\partial^{2}H(x,y;t)}{\partial t^{2}}-\sigma t\triangledown^{2}H(x,y;t)=-\frac{\epsilon_{0}}{2}\Bigl{[}\frac{(V_{\rm bg}-V_{0})^{2}}{H^{2}(x,y;t)}+\frac{V^{2}_{\rm rms}(H^{2}(x,y;t))}{H^{2}(x,y;t)}\Bigl{]}+P_{\rm Cas}(H(x,y;t))~{}.$ (S.5)
@@
$\rho t\frac{\partial^{2}U(x,y;t)}{\partial t^{2}}-\sigma t\triangledown^{2}U(x,y;t)-k_{\rm eff}^{2}U(x,y;t)=-\frac{\epsilon_{0}}{2}\Bigl{[}\frac{(V_{\rm bg}-V_{0})^{2}}{d^{2}}+\frac{V_{\rm rms}^{2}(d)}{d^{2}}\Bigl{]}+P_{\rm Cas}(d)~{},$ (S.6)
@@
$-\sigma t\triangledown^{2}U_{0}(x,y)-k_{\rm eff}^{2}U_{0}(x,y)+\frac{\epsilon_{0}}{2}\Bigl{[}\frac{(V_{\rm bg}-V_{0})^{2}}{d^{2}}+\frac{V_{\rm rms}^{2}(d)}{d^{2}}\Bigl{]}-P_{\rm Cas}(d)=0~{},$ (S.8)
@@
$\rho t\frac{\partial^{2}U_{f}(x,y;t)}{\partial t^{2}}-\sigma t\triangledown^{2}\Bigl{(}U_{0}(x,y)+U_{f}(x,y;t)\Bigl{)}-k_{\rm eff}^{2}\Bigl{(}U_{0}(x,y)+U_{f}(x,y;t)\Bigl{)}+\frac{\epsilon_{0}}{2}\Bigl{[}\frac{(V_{\rm bg}-V_{0})^{2}}{d^{2}}+\frac{V_{\rm rms}^{2}(d)}{d^{2}}\Bigl{]}-P_{\rm Cas}(d)=0~{}.$ (S.9)
By putting together Equations S.8 and S.9, the EoM is reduced into
$\rho t\frac{\partial^{2}U_{f}(x,y;t)}{\partial t^{2}}-\sigma t\triangledown^{2}U_{f}(x,y;t)-k_{\rm eff}^{2}U_{f}(x,y;t)=0~{}.$ (S.10)
@@
$\iint_{A}\Bigl{(}\rho t\frac{\partial^{2}U_{f}(x,y;t)}{\partial t^{2}}-\sigma t\triangledown^{2}U_{f}(x,y;t)-k_{\rm eff}^{2}U_{f}(x,y;t)\Bigl{)}\phi_{1,1}dxdy=0,$ (S.12)
@@
[…]
@@
$\omega_{0}^{2}=2\frac{\pi^{2}}{L^{2}}\frac{\sigma}{\rho}-\frac{k^{2}_{\rm eff}}{\rho t}~{}.$ (S.14)
@@
$\omega_{0}^{2}=\omega_{n}^{2}\Bigl{(}1-\frac{L^{2}k^{2}_{\rm eff}}{2\pi^{2}\sigma t}\Bigl{)}~{}.$ (S.15)
For clarity, we denote temperature-dependent variations in frequencies, pressures, and similar parameters by $\Delta$, while $\delta$ indicates changes at a fixed temperature $T$. From Equation S.14, we can calculate the membrane’s frequency response under a given external force to be
$\delta\omega^{2}=\omega_{0}^{2}-\omega_{{}_{n}}^{2}=-\frac{k^{2}_{\rm eff}}{\rho t}~{},$ (S.16)
@@
$\delta\omega^{2}=-\frac{1}{\rho t}P^{\prime}_{\rm ext}(d)=-\frac{1}{\rho tL^{2}}F^{\prime}_{\rm ext}(d)~{},$ (S.17)
@@
$\Delta\omega^{2}(T_{1},T_{2})=\delta\omega^{2}(T_{2})-\delta\omega^{2}(T_{1})=-\frac{1}{\rho tL^{2}}\bigl{(}F^{\prime}_{\rm ext}(T_{2})-F^{\prime}_{\rm ext}(T_{1})\bigl{)}=-\frac{1}{\rho tL^{2}}\Delta F^{\prime}_{\rm ext}~{}.$ (S.18)
@@
$\frac{\delta\omega}{\omega_{0}}=\frac{\delta f}{f}\approx-\frac{L^{2}}{4\pi^{2}\sigma t}P^{\prime}_{\rm Cas}=-\frac{1}{4\pi^{2}\sigma t}F^{\prime}_{\rm Cas}~{}.$ (S.19)
@@
$\frac{\Delta f}{f}=\frac{\delta f(T_{2})-\delta f(T_{1})}{f}=\frac{1}{4\pi^{2}\sigma t}\Bigl{(}F^{\prime}_{\rm Cas}(T_{2})-F^{\prime}_{\rm Cas}(T_{1})\Bigl{)}~{},$ (S.20)
@@
$\omega_{0}^{2}=2\frac{\pi^{2}}{L^{2}}\frac{\sigma}{\rho}+\frac{B\cdot n}{\rho td^{n+1}},$ (S.22)
@@
$\delta\omega_{0}^{2}=-\frac{P^{\prime}(d)}{\rho t}=-\frac{F^{\prime}(d)}{\rho tL^{2}}~{}.$ (S.23)
@@
where $f_{\theta}=\omega/(2\pi)$ is the linear resonant frequency, $z_{\theta}$ is the center displacement due to the external force, $\delta f_{\theta}$ is the variation in the linear frequency ($\theta=h,nh$), and $Y_{\rm ratio}$ is the hole-dependent ratio. In our case, the radius of the holes is 0.75 $\mu$m and the center-to-center separation is 5.5 $\mu$m, $Y_{\rm ratio}=0.923$ is obtained from finite element method (FEM) simulation.
@@
In order to find out the thermal expansion coefficients of the nanomembranes suspended over the small- and big-gap cavity, as shown in Fig. 5(f) in the main text, we derive the membrane frequency dependence on the temperature here. For squared tensile nanomembranes, the linear resonant frequency $f=\omega/(2\pi)$ of the fundamental mode during the temperature ($T$) sweep is
@@
[…]
@@
With the temperature data points shown in Fig. 5(f) in the main text, one can calculate the thermal expansion coefficients of membranes suspending over the small- and big-gap cavities using Equation S.33. For the small-gap membrane
@@
In this paper, the nanomembrane suspended over a 190 nm gap size experiences a variation of resonant frequency that is different from the variation of the big-gap membrane, in order to find the corresponding abrupt pressure/force change on the small-gap membrane near $T_{C}$, we first perform linear fits on temperature-dependent frequency data points ($\omega=2\pi f$) below $T_{C}$ for both small-gap membrane (Fig. S19d) and big-gap membranes (Fig. S20(e) and Fig. S20(f)), in order to accurately cancel out the thermal expansion of the membranes. Linear fits are used, because for such small temperature span the thermal expansion coefficients can be approximated as constant values. Then we subtract the different between the frequency square values $\omega^{2}$ at a temperature above $T_{C}$ from the linear fit, and obtain the variation of angular frequency square $\Delta\omega^{2}$ for membranes on both gap sizes, which is shown in Fig. S20(a). Later we compare the $\Delta\omega^{2}$ difference between the two membranes, and use Equation S.17 to check the corresponding force gradient and pressure gradient variations, as shown in Fig. S20(b) and S20(c), respectively, without knowing the exactly behaviors of the external pressure/force. The error bars in Fig. S20(a) are obtained with the different frequency sweep data points for every temperature setting, while the error bars in Fig. S20(b) and S20(c) are calculated by adding up the error bars of $\Delta\omega^{2}$ for small- and big-gap membranes shown in Fig. S20(a). Here we combined the datasets from the big-gap membranes into one dataset. The error bars of the $\Delta\omega^{2}$ data on membranes over both gap sizes are obtained from the frequency sweep datasets at each temperature data point, and the error bars of the force gradient and pressure gradient variations are calculated by adding up the error bar size of the small-gap membrane at a certain temperature and the two error bar sizes of the big-gap membrane around the corresponding temperature.
Figure S20: Procedure on calculating the pressure and force variations from experimental frequency shifts.After performing linear fits on the variation of $\omega^{2}$ for the small-gap and big-gap membranes, the derivations of $\omega^{2}$ from the linear fits are shown in Panel (a) for small- and big-gap membranes, where the two big-gap membrane datasets are combine. After subtracting the difference in $\omega^{2}$ between small- and big-gap membranes, we used Equation S.17 to obtain the corresponding force gradient and pressure gradient variations, as shown in Panels (b) and (c). Here the exact form of the external force/pressure is not required to be known beforehand for the calculation.
@@
Figure S21: Estimation of the Casimir pressure between superconductors, at the temperature of $T=0.99\cdot T_{C}$. (a) Casimir pressure between superconducting NbTiN surfaces. (b) Power law value $x$ in the Casimir pressure $P(d)\sim 1/d^{n}$ between superconducting NbTiN surfaces. (c) Casimir pressure between superconducting NbTiN surfaces, with a separation around 190 nm. (d) Power law value $x$ in the Casimir pressure $P(d)~{}d^{n}$ between superconducting NbTiN around $T_{C}$, around 190 nm separation. At 190 nm, the absolute Casimir pressure between NbTiN near $T_{C}$, is $P=-0.4021$ Pa, and the power law value is $n=3.5069$, leading to the explicit expression of the Casimir pressure to be $P_{\rm Cas}\bigl{|}_{d=190\rm nm}=(-1.081\times 10^{-24})/d^{3.507}$ [Pa].
@@
[…]
@@
where $\Omega$ represents the plasma frequency for intraband transitions, and $\gamma$ is the relaxation frequency. We have assumed the plasma frequency, $\Omega$, to be temperature-independent, using its room-temperature value. The relaxation frequency, $\gamma$, on the other hand, varies with temperature, typically decreasing as temperature decreases. At cryogenic temperatures, $\gamma$ approaches a constant residual value that depends on the specific sample. Following standard practice, we relate the residual relaxation frequency to the room-temperature value, $\gamma_{0}$, using the formula $\gamma=\gamma_{0}/\rm RRR$, where RRR is the residual resistance ratio. The following Drude parameters were used: $\Omega=5.33$ $\hbar$ and $\gamma_{0}=0.465$ $\hbar$. Moreover, we took $\text{RRR}=1$. The permittivity of NbTiN in the superconducting state was calculated using the Mattis-Bardeen formula for the conductivity $\sigma_{\rm BCS}(\omega)$ [105], known to accurately represent the optical response of BCS superconductors [106]. While the general Mattis-Bardeen formula depends on both frequency ($\omega$) and wavevector ($q$), the $q$-dependence is negligible in the dirty limit ($\ell/\xi_{0}\ll 1$), where $\ell=v_{F}/\gamma$ the dirty limit is well satisfied ($\ell/\xi_{0}=1.4\times 10^{-2}$). This is confirmed by optical measurements of NbTiN films in the THz region, that are in excellent agreement with the local dirty-limit of the Mattis-Bardeen formula [107]. Analytic continuation to imaginary frequencies [108], gives:
@@
where $\Omega$ is the plasma frequency for intraband transition, and $\gamma$ is the relaxation frequency. For NbTiN we adapt the values used from [14, 17], i.e. Ω=5.33eV/$\Omega=5.33~{}\text{eV}/\hbar$ and $\gamma=0.465~{}\text{eV}/\hbar$. The function $g(\xi;T)$ that describes the modification as a result of BCS theory, can be expressed as
@@
From data shown in Fig. S21 the effective Casimir pressure between NbTiN films at a separation of 190 nm, near $T_{C}$, can be expressed as
@@
with the separation $d$ measured in meters. With this expression of the Casimir pressure, we can use FEM simulation to estimate how much the external Casimir pressure/force is acting on the membrane. In order to do that, we scale the Casimir pressure acting on the membrane by a factor $\beta$ close to unity, i.e. $P_{\rm sim}=(1+\beta)\cdot P_{\rm Cas}(d)$ with $|\beta|<0.001$, and found the linear conversion factors from the the frequency shift $\Delta\omega^{2}$ to variations in the Casimir force $\Delta F_{\rm Cas}$, the Casimir pressure $\Delta P_{\rm Cas}$, and membrane center static deflection $\Delta z_{\rm Cas}$. The conversion factors are $1.986\times 10^{-17}~{}\rm N/Hz^{2}$, $3.941\times 10^{-11}~{}\rm Pa/Hz^{2}$, and $1.592\times 10^{-20}~{}\rm m/Hz^{2}$, for force, pressure, and center deflection, respectively. From the simulation results, one can calculate that for the membrane suspends over 190 nm gap, the total Casimir force is around $-2.028\times 10^{-7}$ N, and the total center deflection is around $-141.6$ pm, indicating that the level of parallelism achieved in our device is state-of-the-art.
@@
In this Section we study the $n=0$ contribution to the Lifshitz formula  S.38. This term represents a classical contribution to the Casimir pressure, which becomes the dominant contribution for separations $d\gg\lambda_{T}$, where $\lambda_{T}=\hbar c/k_{B}T$ is the thermal length. The precise form of this contribution has been a matter of significant controversy within the Casimir community for more than two decades, with experts divided between the Drude and plasma models. The Drude model accounts for electron scattering and dissipation, offering a more intuitive physical interpretation, while the plasma model, despite neglecting dissipation, aligns more closely with experimental observations. The key divergence lies in the treatment of the zero-frequency Matsubara term in the Lifshitz formula: the Drude approach predicts a vanishing transverse electric (TE) contribution, whereas the plasma model yields a finite term that significantly influences the force magnitude. This debate not only reflects discrepancies between theory and experiment but also challenges our fundamental understanding of the interplay between dissipation and quantum fluctuations.
The zero-frequency ($n=0$) term in the discrete Matsubara frequency sum represents a static contribution to Lifshitz formulae Equations S.38 and S.41. It is well-established that this single term plays a dominant role in determining the temperature dependence of the Casimir pressure [112, 113, 114]. The $n=0$ term can be separated into TM and TE polarization contributions:
@@
[…]
@@
[…]
@@
[…]
@@
Figure S23: Theoretical predictions for the Casimir pressure gradient variation $\Delta P^{\prime}$ and the associated membrane frequency square shift $\Delta\omega^{2}$ across the superconducting transition. Panels (a) and (b) show curves computed using BCS model in the superconducting state ($T<T_{C}$) and the Drude model in the normal state ($T>T_{C}$). Panels (c) and (d) show curves computed using BCS model in the superconducting state ($T<T_{C}$) and the plasma model in the normal state.($T>T_{C}$). Panels (e) and (f) show curves computed using plamsa model both in the superconducting state ($T<T_{C}$) and in the normal state.($T>T_{C}$).

### v2

@@ Acknowledgements.
The resonance frequency of the parallel plate capacitor, consists of a suspended NbTiN membrane under the electrostatic and the Casimir interaction with the fixed NbTiN backgate, is calculation in this Supplementary Information. The schematic of the device is illustrated in Figure S19. We focus on analyzing the fundamental mode frequency $\omega_{0}$ (or $\omega_{\rm(1,1)}$). The thin membrane is under high tension, and the height of the membrane is deviated from the equilibrium (not forced) position only by a negligible amount compared to the separation. Based on these considerations, the membrane’s dynamics are described by the following equation of motion (EoM) [53]:
$\rho h\frac{\partial^{2}H(x,y;t)}{\partial t^{2}}-\sigma h\triangledown^{2}H(x,y;t)=P_{\rm ES}(H(x,y;t))+P_{\rm Cas}(H(x,y;t))~,$ (S.8)
where the height function $H(x,y;t)$ describes the separation between the membrane and the backgate, $\rho$ the density, $\sigma$ the tensile stress, and $h$ the thickness of the membrane. $P_{\rm ES}(H(x,y;t))$ and $P_{\rm Cas}(H(x,y;t))$ are the pressures coming from the electrostatic force and the Casimir force between the membrane and the backgate, respectively. The electrostatic and Casimir forces can be estimated using the additive Derjaguin Proximity Force Approximation (PFA) [54]. This method decomposes the NbTiN surfaces into parallel patch pairs and sums the corresponding pairwise force contributions, thus
$\rho h\frac{\partial^{2}H(x,y;t)}{\partial t^{2}}-\sigma h\triangledown^{2}H(x,y;t)=-\frac{\epsilon_{0}}{2}\Bigl[\frac{(V_{\rm bg}-V_{0})^{2}}{H^{2}(x,y;t)}+\frac{V^{2}_{\rm rms}(H^{2}(x,y;t))}{H^{2}(x,y;t)}\Bigl]+P_{\rm Cas}(H(x,y;t))~.$ (S.9)
@@
$\rho h\frac{\partial^{2}U(x,y;t)}{\partial t^{2}}-\sigma h\triangledown^{2}U(x,y;t)-k_{\rm eff}^{2}U(x,y;t)=-\frac{\epsilon_{0}}{2}\Bigl[\frac{(V_{\rm bg}-V_{0})^{2}}{d^{2}}+\frac{V_{\rm rms}^{2}(d)}{d^{2}}\Bigl]+P_{\rm Cas}(d)~,$ (S.10)
@@
$\sigma h\triangledown^{2}U_{0}(x,y)-k_{\rm eff}^{2}U_{0}(x,y)+\frac{\epsilon_{0}}{2}\Bigl[\frac{(V_{\rm bg}-V_{0})^{2}}{d^{2}}+\frac{V_{\rm rms}^{2}(d)}{d^{2}}\Bigl]-P_{\rm Cas}(d)=0~,$ (S.12)
@@
$\rho h\frac{\partial^{2}U_{f}(x,y;t)}{\partial t^{2}}-\sigma h\triangledown^{2}\Bigl(U_{0}(x,y)+U_{f}(x,y;t)\Bigl)-k_{\rm eff}^{2}\Bigl(U_{0}(x,y)+U_{f}(x,y;t)\Bigl)+\frac{\epsilon_{0}}{2}\Bigl[\frac{(V_{\rm bg}-V_{0})^{2}}{d^{2}}+\frac{V_{\rm rms}^{2}(d)}{d^{2}}\Bigl]-P_{\rm Cas}(d)=0~.$ (S.13)
By putting together Equations S.12 and S.13, the EoM is reduced into
$\rho h\frac{\partial^{2}U_{f}(x,y;t)}{\partial t^{2}}-\sigma h\triangledown^{2}U_{f}(x,y;t)-k_{\rm eff}^{2}U_{f}(x,y;t)=0~.$ (S.14)
@@
$\iint_{A}\Bigl(\rho h\frac{\partial^{2}U_{f}(x,y;t)}{\partial t^{2}}-\sigma h\triangledown^{2}U_{f}(x,y;t)-k_{\rm eff}^{2}U_{f}(x,y;t)\Bigl)\phi_{1,1}dxdy=0,$ (S.16)
@@
$\rho h\omega_{0}^{2}\iint_{A}\phi_{1,1}^{2}dxdy+2\frac{\pi^{2}}{L^{2}}\sigma h\iint_{A}\phi_{1,1}^{2}dxdy-k^{2}_{\rm eff}\iint_{A}\phi_{1,1}^{2}dxdy=0.$ (S.17)
@@
$\omega_{0}^{2}=2\frac{\pi^{2}}{L^{2}}\frac{\sigma}{\rho}-\frac{k^{2}_{\rm eff}}{\rho h}~.$ (S.18)
@@
$\omega_{0}^{2}=\omega_{n}^{2}\Bigl(1-\frac{L^{2}k^{2}_{\rm eff}}{2\pi^{2}\sigma h}\Bigl)~.$ (S.19)
From Equation S.18, we can calculate the membrane’s frequency response under a given external force to be (For clarity, we denote temperature-dependent variations in frequencies, pressures, and similar parameters by $\Delta$, while $\delta$ indicates changes at a fixed temperature $T$.)
$\delta\omega^{2}=\omega_{0}^{2}-\omega_{{}_{n}}^{2}=-\frac{k^{2}_{\rm eff}}{\rho h}~,$ (S.20)
@@
$\delta\omega^{2}=-\frac{1}{\rho h}P^{\prime}_{\rm ext}(d)=-\frac{1}{\rho hL^{2}}F^{\prime}_{\rm ext}(d)~,$ (S.21)
@@
$\Delta\omega^{2}(T_{1},T_{2})=\delta\omega^{2}(T_{2})-\delta\omega^{2}(T_{1})=-\frac{1}{\rho hL^{2}}\bigl(F^{\prime}_{\rm ext}(T_{2})-F^{\prime}_{\rm ext}(T_{1})\bigl)=-\frac{1}{\rho hL^{2}}\Delta F^{\prime}_{\rm ext}~.$ (S.22)
@@
$\frac{\delta\omega}{\omega_{0}}=\frac{\delta f}{f}\approx-\frac{L^{2}}{4\pi^{2}\sigma h}P^{\prime}_{\rm Cas}=-\frac{1}{4\pi^{2}\sigma h}F^{\prime}_{\rm Cas}~.$ (S.23)
@@
$\frac{\Delta f}{f}=\frac{\delta f(T_{2})-\delta f(T_{1})}{f}=\frac{1}{4\pi^{2}\sigma h}\Bigl(F^{\prime}_{\rm Cas}(T_{2})-F^{\prime}_{\rm Cas}(T_{1})\Bigl)~,$ (S.24)
@@
$\omega_{0}^{2}=2\frac{\pi^{2}}{L^{2}}\frac{\sigma}{\rho}+\frac{B\cdot n}{\rho hd^{n+1}},$ (S.26)
@@
$\delta\omega_{0}^{2}=-\frac{P^{\prime}(d)}{\rho h}=-\frac{F^{\prime}(d)}{\rho hL^{2}}~.$ (S.27)
@@
[…]
@@
In order to find out the thermal expansion coefficients of the nanomembranes suspended over the small- and big-gap cavity, as shown in Figure 5(f) in the main text, we derive the membrane frequency dependence on the temperature here. For squared tensile nanomembranes, the linear resonant frequency $f=\omega/(2\pi)$ of the fundamental mode during the temperature ($T$) sweep is
@@
$\sigma_{\text{th}}=\frac{E_{\text{film}}}{1-\nu_{\text{film}}}\int^{T_{2}}_{T_{1}}(\alpha_{\text{film}}(T)-\alpha_{\text{sub}}(T))dT,$ (S.38)
@@
With the temperature data points shown in Figure 5(f) in the main text, one can calculate the thermal expansion coefficients of membranes suspending over the small- and big-gap cavities using Equation S.37. For the small-gap membrane
@@
Supplementary Information F4: Frequency noise in resonance frequency measurement
For a mechanical resonator driven at an angular frequency $\omega=2\pi f$, the steady-state response $H(\omega)$ is proportional to
$H(\omega)=\frac{1}{\omega_{0}^{2}-\omega^{2}+i\frac{\omega\omega_{0}}{Q}},$ (S.41)
where $\omega_{0}=2\pi f_{0}$. Let $y=\omega/\omega_{0}=f/f_{0}$. The phase lag of the response is
$\phi(f)=\arg(H(\omega))=-\arctan(\frac{\omega\omega_{0}/Q}{\omega_{0}^{2}-\omega^{2}})=-\arctan(\frac{y/Q}{1-y^{2}}).$ (S.42)
Use $\arctan(u)+\arctan(1/u)=\pi/2$ for $u>0$,
$\phi(f)=\frac{\pi}{2}-\arctan[Q(\frac{1}{y}-y)]=\frac{\pi}{2}-\arctan[Q(\frac{f_{0}}{f}-\frac{f}{f_{0}})].$ (S.43)
When $|f-f_{0}|\ll f_{0}$, one has
$\frac{d\phi}{df}=-\frac{Q}{1+[Q(\frac{f_{0}}{f}-\frac{f}{f_{0}})]^{2}}\cdot(-\frac{f_{0}}{f^{2}}-\frac{1}{f_{0}}),$ (S.44)
and
$\frac{d\phi}{df}|_{f=f_{0}}=\frac{2Q}{f_{0}}.$ (S.45)
Thus
$\delta f\approx\frac{f_{0}}{2Q}\delta\phi.$ (S.46)
A lock-in measuring the signal $S_{T}$ with the white noise $N_{T}$ over a bandwidth $B_{T}$, yields a rms phase uncertainty [27]
$\sigma_{\phi}\approx\frac{N_{T}}{S_{T}}\sqrt{B_{T}}=\frac{N_{T}}{S_{T}}\sqrt{\frac{1}{2\pi\tau}},$ (S.47)
here $\tau$ is the integrating time. Converting the phase noise into the frequency noise gives
$\langle\delta f\rangle\approx\frac{f_{0}}{2Q}\sigma_{\phi}=\frac{f_{0}}{2Q}\frac{N_{T}}{S_{T}}\sqrt{\frac{1}{2\pi\tau}}.$ (S.48)
@@
In this paper, the nanomembrane suspended over a 190 nm gap size experiences a variation of resonant frequency that is different from the variation of the big-gap membrane, in order to find the corresponding abrupt pressure/force change on the small-gap membrane near $T_{C}$, we first perform linear fits on temperature-dependent frequency data points ($\omega=2\pi f$) below $T_{C}$ for both small-gap membrane (Figure S21d) and big-gap membranes (Figure S21(e) and Figure S21(f)), in order to accurately cancel out the thermal expansion of the membranes. Linear fits are used, because for such small temperature span the thermal expansion coefficients can be approximated as constant values. Then we subtract the different between the frequency square values $\omega^{2}$ at a temperature above $T_{C}$ from the linear fit, and obtain the variation of angular frequency square $\Delta\omega^{2}$ and linear frequency $\Delta f$ for membranes on both gap sizes, which is shown in Figure S22(a) and Figure S22(b), respectively. Later we compare the $\Delta\omega^{2}$ difference between the two membranes, and use Equation S.21 to check the corresponding force gradient and pressure gradient variations, as shown in Figure S22(c) and S22(d), respectively, without knowing the exactly behaviors of the external pressure/force. The error bars in Figure S22(a) and Figure S22(b) are obtained with the different frequency sweep data points for every temperature setting, while the error bars in Figure S22(c) and S22(d) are calculated by adding up the error bars of $\Delta\omega^{2}$ for small- and big-gap membranes shown in Figure S22(a). Here we combined the datasets from the big-gap membranes into one dataset. The error bars of the $\Delta\omega^{2}$ data on membranes over both gap sizes are obtained from the frequency sweep datasets at each temperature data point, and the error bars of the force gradient and pressure gradient variations are calculated by adding up the error bar size of the small-gap membrane at a certain temperature and the two error bar sizes of the big-gap membrane around the corresponding temperature.
Figure S22: Procedure on calculating the pressure and force variations from experimental frequency shifts.After performing linear fits on the variation of $\omega^{2}$ for the small-gap and big-gap membranes, the derivations of $\omega^{2}$ from the linear fits are shown in Panel (a) for small- and big-gap membranes, where the two big-gap membrane datasets are combine. The derivations of $f$ from the linear fits are shown in Panel (b). After subtracting the difference in $\omega^{2}$ between small- and big-gap membranes, we used Equation S.21 to obtain the corresponding force gradient and pressure gradient variations, as shown in Panels (c) and (d). Here the exact form of the external force/pressure is not required to be known beforehand for the calculation.
@@
Figure S23: Estimation of the Casimir pressure between superconductors, at the temperature of $T=0.99\cdot T_{C}$. (a) Casimir pressure between superconducting NbTiN surfaces. (b) Power law value $x$ in the Casimir pressure $P(d)\sim 1/d^{n}$ between superconducting NbTiN surfaces. (c) Casimir pressure between superconducting NbTiN surfaces, with a separation around 190 nm. (d) Power law value $x$ in the Casimir pressure $P(d)~d^{n}$ between superconducting NbTiN around $T_{C}$, around 190 nm separation. At 190 nm, the absolute Casimir pressure between NbTiN near $T_{C}$, is P=0.4021P=-0.4021 Pa, and the power law value is $n=3.5069$, leading to the explicit expression of the Casimir pressure to be $P_{\rm Cas}\bigl|_{d=190\rm nm}=(-1.081\times 10^{-24})/d^{3.507}$ [Pa].
@@
$P^{\prime}(d,T)=2\frac{k_{B}T}{\pi}\sum^{\infty}_{l=0}{}^{\prime}\int^{\infty}_{0}dk_{\bot}k_{\bot}q_{l}^{2}\sum_{\alpha}\frac{e^{2dq_{l}}r_{\alpha}^{2}(i\xi_{l},k_{\bot})}{[e^{2dq_{l}}-r_{\alpha}^{2}(i\xi_{l},k_{\bot})]^{2}},$ (S.53)
@@
[…]
@@
where $\Omega$ is the plasma frequency for intraband transition, and $\gamma$ is the relaxation frequency. For NbTiN we adapt the values used from [14, 16], i.e. $\Omega=5.33~\text{eV}/\hbar$ and $\gamma=0.465~\text{eV}/\hbar$. The function $g(\xi;T)$ that describes the modification as a result of BCS theory, can be expressed as
@@
From data shown in Figure S23 the effective Casimir pressure between NbTiN films at a separation of 190 nm, near $T_{C}$, can be expressed as
@@
with the separation $d$ measured in meters. With this expression of the Casimir pressure, we can use FEM simulation to estimate how much the external Casimir pressure/force is acting on the membrane. In order to do that, we scale the Casimir pressure acting on the membrane by a factor $\beta$ close to unity, i.e. $P_{\rm sim}=(1+\beta)\cdot P_{\rm Cas}(d)$ with $|\beta|<0.001$, and found the linear conversion factors from the the frequency shift $\Delta\omega^{2}$ to variations in the Casimir force $\Delta F_{\rm Cas}$, the Casimir pressure $\Delta P_{\rm Cas}$, and membrane center static deflection $\Delta z_{\rm Cas}$. The conversion factors are $7.83\times 10^{-16}~\rm N/Hz^{2}$, $1.55\times 10^{-9}~\rm Pa/Hz^{2}$, and $6.28\times 10^{-19}~\rm m/Hz^{2}$, for force, pressure, and center deflection, respectively. From the simulation results, one can calculate that for the membrane suspends over 190 nm gap, the total Casimir force is around $2.028\times 10^{-7}$ N, and the total center deflection is around 153-153 pm, indicating that the level of parallelism achieved in our device is state-of-the-art.
@@
In this Section we study the $n=0$ contribution to the Lifshitz formula  S.50. This term represents a classical contribution to the Casimir pressure, which becomes the dominant contribution for separations $d\gg\lambda_{T}$, where $\lambda_{T}=\hbar c/k_{B}T$ is the thermal length. The precise form of this contribution has been a matter of significant controversy within the Casimir community for more than two decades, with experts divided between the Drude and plasma models. The Drude model accounts for electron scattering and dissipation, offering a more intuitive physical interpretation, while the plasma model, despite neglecting dissipation, aligns more closely with experimental observations. The key divergence lies in the treatment of the zero-frequency Matsubara term in the Lifshitz formula: the Drude approach predicts a vanishing transverse electric (TE) contribution, whereas the plasma model yields a finite term that significantly influences the force magnitude. This debate not only reflects discrepancies between theory and experiment but also challenges our fundamental understanding of the interplay between dissipation and quantum fluctuations.
The zero-frequency ($n=0$) term in the discrete Matsubara frequency sum represents a static contribution to Lifshitz formulae Equations S.50 and S.53. It is well-established that this single term plays a dominant role in determining the temperature dependence of the Casimir pressure [65, 66, 67]. The $n=0$ term can be separated into TM and TE polarization contributions:
@@
$r_{\text{TE}}(0,k_{\bot})=0,\quad\quad\text{(Drude model approach)}~.$ (S.66)
@@
$P_{\rm TE}^{(0)}=0,\quad\quad\text{(Drude model approach)}~.$ (S.67)
@@
$r_{\rm TE}(0,k_{\bot})=\frac{k_{\bot}-\sqrt{\Omega^{2}/c^{2}+k_{\bot}^{2}}}{k_{\bot}+\sqrt{\Omega^{2}/c^{2}+k_{\bot}^{2}}},\quad\quad\text{(plasma model approach)}~.$ (S.68)
@@
Figure S25: Theoretical predictions for the Casimir pressure gradient variation $\Delta P^{\prime}$ and the associated membrane frequency square shift $\Delta\omega^{2}$ across the superconducting transition. Panels (a), (b) and (c) show curves computed using BCS model in the superconducting state ($T<T_{C}$) and the Drude model in the normal state ($T>T_{C}$). Panels (d), (e) and (f) show curves computed using BCS model in the superconducting state ($T<T_{C}$) and the plasma model in the normal state.($T>T_{C}$). Panels (g), (h) and (I) show curves computed using plasma model both in the superconducting state ($T<T_{C}$) and in the normal state.($T>T_{C}$).